OpenStudy (anonymous):

How do you factor polynomials with cubic powers and polynomials with powers higher than cubic? help please.

OpenStudy (kc_kennylau):

the simple answer is "no you don't"

OpenStudy (kc_kennylau):

the more complicated answer is you can try factors by the factor theorem

OpenStudy (kc_kennylau):

and the factors are often related to the constant term, e.g. the constant term is 9 and the factor can be x-9 or x-3 or x-1 or x-1/3 or x-1/9 or x+9 or x+3 or x+1 or x+1/3 or x+1/9 etc

OpenStudy (kc_kennylau):

If he says that the factors are integral you can try x=+-1 or x=+-2 or x=+-3 and don't forget x=0

OpenStudy (kc_kennylau):

in short, don't mess up with 'em hope i helped you :)

OpenStudy (anonymous):

i wish it did but i have to know how to use them with synthetic division on my algebra two quiz and i know how but the whole p and q rules that go with it are kind of confusion

OpenStudy (kc_kennylau):

do you have an example? :)

OpenStudy (anonymous):

3x^4- 75x^2 or 1-27x^6 either or will help just fine

OpenStudy (kc_kennylau):

oh only two terms

OpenStudy (kc_kennylau):

\[3x^4-75x^2\]\[=x^2(3x^2-75)\]\[=3x^2(x^2-25)\]\[=3x^2(x+5)(x-5)\] \[1-27x^6\]\[=(1-3x)(1+3x+9x^2)\]

OpenStudy (anonymous):

thank you so much

OpenStudy (kc_kennylau):

no problem :) just memorize the 5 identities by heart :)

OpenStudy (kc_kennylau):

1 and 2 \((x\pm y)^2=x^2\pm2xy+y^2\) 3 and 4 \(x^3\pm y^3=(x\pm y)(x^2\mp xy+y^2)\) 5 \(x^2-y^2=(x+y)(x-y)\)